# Certifiably Optimal Anisotropic Rotation Averaging

### v1

@@ Abstract
Rotation averaging is a key subproblem in applications of computer vision and robotics. Many methods for solving this problem exist, and there are also several theoretical results analyzing difficulty and optimality. However, one aspect that most of these have in common is a focus on the isotropic setting, where the intrinsic uncertainties in the measurements are not fully incorporated into the resulting optimization task. Recent empirical results suggest that moving to an anisotropic framework, where these uncertainties are explicitly included, can result in an improvement of solution quality. However, global optimization for rotation averaging has remained a challenge in this scenario. In this paper we show how anisotropic costs can be incorporated in certifiably optimal rotation averaging. We also demonstrate how existing solvers, designed for isotropic situations, fail in the anisotropic setting. Finally, we propose a stronger relaxation and show empirically that it is able to recover global optima in all tested datasets and leads to a more accurate reconstruction in all but one of the scenes.
@@ 1 Introduction
In large part this is because of the great importance it has for many problems within computer vision and robotics. Traditionally, it has been a core component of non-sequential structure from motion (SfM), see e.g. [26, 27, 32]. The key strength here is that rotation averaging optimizes the orientations of all the cameras at the same time, allowing recovery methods to avoid incrementally increasing errors (drift) [12] that easily occur in sequential SfM methods. More recently, [5, 7, 8] also point out the importance of rotation averaging for improving accuracy and speed of simultaneous localization and mapping (SLAM).
@@
Further, we empirically verify—on a number of synthetic and real datasets—that this new relaxation is sufficient to recover the globally optimal solution, and that our anisotropic model generally achieves solutions that are more accurate than the standard isotropic chordal penalty.
In summary, our main contributions are:
@@ 2 Related Work
The works [2, 21] also considers certifiable optimization of the anisotropic chordal distance. However, their SDP relaxation is based on the Cayley mapping and the relation to our formulation is somewhat unclear. In addition, [2, 21] have to introduce redundant constraints in order to get a “reasonably tight” SDP relaxation. It is likely, in the light of our analysis in this paper, and the discussion in [3] that this is a result of the authors only enforcing $O(3)$ membership for their optimization.
It is also worth mentioning that—while this paper looks at the anisotropic $L_{2}$ chordal distance—, it is possible to consider other distances between rotations as well as their robust counterparts. E.g., [43, 44] consider the geodesic distance (i.e. the one on the manifold of $SO(3)$). Further, in [9, 39, 42] the authors analyze and look at different (robustified) distances in order to obtain a “robust” formulation (resulting in the solution being less susceptible to outliers). Outliers are handled via the introduction of auxiliary variables in the SDP formulation in [35, 6].
From an algorithmic point of view, the SPD relaxation can be solved by general-purpose solvers, e.g. [1]. There are also several dedicated solvers, constructed specifically for the rotation averaging problem. [16, 33] use coordinate descent (fixing all but one row/column and minimizing), [25] use a primal-dual update rule, and in [10, 45] the authors combine a global coordinate descent approach with a local step. An early approach [40] looks at eigenvectors of a specific matrix. A particularly powerful method used in [4, 14] is that of the Riemannian staircase, where one optimizes using block matrices of a certain dimension and increases (“climb the staircase”) the dimension until a global optimum is found. The authors show that such a method can be made highly effective.
@@ 3 Certifiably Optimal Rotation Averaging
The derivation of a strong convex relaxation relies on two observations: Firstly, since the set of rotation matrices have constant (Frobenius) norm, the objective function can be replaced by $-\sum_{i\neq j}\langle\tilde{R}_{ij},R_{j}R_{i}^{\top}\rangle$. If we let $\mathbf{R}=\begin{bmatrix}R_{1}^{\top}&R_{2}^{\top}&\ldots\end{bmatrix}^{\top}$ and $\mathbf{\tilde{R}}$ be the matrix containing the blocks $\tilde{R}_{ij}$ (and zeros where no relative rotation estimate is available) we can simplify the objective to $-\langle\mathbf{\tilde{R}},\mathbf{R}\mathbf{R}^{\top}\rangle$, which is linear in the relative rotations. Secondly, by orthogonality we have $R_{i}R_{i}^{\top}={\mathtt{I}}$, resulting in the Quadratically Constrained Quadratic Program (QCQP)
@@
[…]
The above program is convex and can therefore be solved reliably using general purpose [1, 23] or specialized solvers [14, 4, 16]. The difference $p^{*}-d_{2}^{*}$ between objective values is referred to as the duality gap. It is clear that we can construct $\mathbf{X}=\mathbf{R}\mathbf{R}^{\top}$ from a solution $\mathbf{R}$ of (2), which is feasible in (SDP-O(3)-iso) and therefore $p^{*}\geq d_{2}^{*}$. Moreover, if the solution $\mathbf{X}$ to (SDP-O(3)-iso) has $\operatorname{rank}(\mathbf{X})=3$ then it can be factorized into $\mathbf{X}=\mathbf{R}\mathbf{R}^{\top}$, with $\mathbf{R}$ being feasible in (2), which yields $p^{*}=d_{2}^{*}$ and certifies that $\mathbf{R}$ is globally optimal in (2). Thus, a solution to the original non-convex QCQP is obtained through optimization via convex SDP program.
@@ 4 Anisotropic Rotation Averaging
While the use of the chordal distance often results in tight SDP relaxations, a downside is that it does not take into account the uncertainty of the relative rotation estimates $\tilde{R}_{ij}$. For two-view relative pose problems, it is frequently the case that the reprojection errors are much less affected by rotations in certain directions than others. In Figure 1 we illustrate this on a real two-view problem. The left image shows the ground truth solution (obtained after bundle adjustment). To generate the graphs to the right, we sampled rotations of the second camera around its x-,y-, and z-axes between $-5$ and $5$ degrees. For each sampled rotation, we found the best 3D point locations and position of the second camera, and computed the sum-of-squared (calibrated) reprojection errors (solid curves). The result shows that rotations around the black/upward axis affects the reprojection errors significantly less than the others, resulting in a direction of larger uncertainty. By performing the sensitivity analysis (i.e. by looking at the singular values of the Jacobian) one can arrive at the similar conclusion. It is therefore desirable to propagate this information to the rotation averaging stage to increase the accuracy of the model. The dashed curves (that are almost indistinguishable from the solid ones) show the approximations, presented below, that can be used for this purpose. In the supplementary material, we also show how isotropic rotation averaging gets negatively affected by the single noisy relative rotation (while the proposed method presented further remains unaffected).
@@ 4.1 Incorporating uncertainties
Consider a relative rotation $\tilde{R}$ that has been optimized to a minimum using a nonlinear least squares approach in a local two-view optimization. Using the exponential map we write $R=e^{[\Delta\bm{\omega}]_{\times}}\tilde{R}$ [^1] Other parameterizations are discussed in the supplementary material., where $\Delta\bm{\omega}$ is an axis-angle representation of the deviation from $\tilde{R}$, i.e. the axis-angle of $R\tilde{R}^{\top}$. Since $\tilde{R}$ is a minimum, gradient terms vanish and the two-view objective can locally be approximated by a quadratic function (ignoring an inconsequential constant)
$\tfrac{1}{2}\Delta\bm{\omega}^{\top}H\Delta\bm{\omega}.$ (3)
The positive semidefinite matrix $H=J^{\top}J$ (where $J$ is the Jacobian of the residual function) approximates the growth of the objective function in a region around the minimum $\tilde{R}$. We remark that the initial two-view objective function can depend on additional parameters such as camera positions and 3D point locations. Here we assume that these “nuisance” parameters have been marginalized out and therefore work with a reduced objective solely in terms of the rotation parameters $\Delta\bm{\omega}$ (using the Schur complement to factor out additional unknowns).
We assume Laplace’s approximation to be valid at least locally in order to quantify the uncertainty of $\Delta\bm{\omega}$, i.e. identify the precision matrix with the Hessian $H$. While in general the uncertainty of a $3\times 3$ matrix $R$ is represented by a $9\times 9$ covariance/precision matrix, we aim for a linear cost (in terms of rotation matrices) in order to achieve a strong relaxation, which means that the objective is restricted to an inner product between $R$ and some other (input-dependent) matrix. We first find a matrix $M$ such that
$\begin{split}\Delta\bm{\omega}^{\top}H\Delta\bm{\omega}&=\operatorname{tr}\left([\Delta\bm{\omega}]_{\times}^{\top}M[\Delta\bm{\omega}]_{\times}\right)\\ &=-\operatorname{tr}\left(M[\Delta\bm{\omega}]_{\times}^{2}\right).\end{split}$ (4)
Using the identity $[\Delta\bm{\omega}]_{\times}^{2}=\Delta\bm{\omega}^{\top}\Delta\bm{\omega}{\mathtt{I}}-\Delta\bm{\omega}\Delta\bm{\omega}^{\top}$, we obtain that $H=\operatorname{tr}(M){\mathtt{I}}-M$. Taking the trace on both sides further shows that $\operatorname{tr}(H)=2\operatorname{tr}(M)$, resulting in
$M=\tfrac{\operatorname{tr}(H)}{2}{\mathtt{I}}-H.$ (5)
@@
$[\Delta\bm{\omega}]_{\times}\approx R\tilde{R}^{\top}-{\mathtt{I}},$ (6)
@@
$\tfrac{1}{2}\operatorname{tr}\big{(}(R\tilde{R}^{\top}-{\mathtt{I}})^{\top}M(R\tilde{R}^{\top}-{\mathtt{I}})\big{)}.$ (7)
@@
$-\operatorname{tr}\big{(}M\tilde{R}R^{\top}\big{)}=-\langle M\tilde{R},R\rangle.$ (8)
We observe that this objective can be viewed as a negative log-likelihood of a Langevin distribution $\mathcal{L}(R,M)$, e.g. [11], with a density function of the type
$p(\tilde{R})\propto e^{\langle M\tilde{R},R\rangle},$ (9)
and that generating rotations $\tilde{R}$ according to (9) is approximately (up to second order terms) equivalent to generating axis-angle vectors $\Delta\bm{\omega}$ from a normal distribution $\mathcal{N}(0,H^{-1})$ with density
$p(\Delta\bm{\omega})\propto e^{-\frac{1}{2}\Delta\bm{\omega}^{\top}H\Delta\bm{\omega}}.$ (10)
Hence, the above derivations suggest a way of generalizing the rotation averaging approach by replacing the relative estimates $\tilde{R}_{ij}$ with a weighted version $M_{ij}\tilde{R}_{ij}$, where $M_{ij}$ is computed from the Hessian $H_{ij}$ of the relative pose problem as outlined above. Therefore, our proposed objective to minimize is
$-\sum\nolimits_{i\neq j}\langle M_{ij}\tilde{R}_{ij},R_{j}R_{i}^{\top}\rangle=-\langle\mathbf{N},\mathbf{R}\mathbf{R}^{\top}\rangle,$ (11)
where $\mathbf{N}$ is a symmetric block matrix containing the blocks $(M_{ij}\tilde{R}_{ij})^{\top}$ for $i<j$, 00 when $i=j$, and $M_{ji}\tilde{R}_{ji}$ for $i>j$.
We conclude this section by remarking that—in contrast to $H$—the matrix $M$ will generally not be positive semidefinite. The following lemma characterizes its eigenvalues
@@
If $H\succeq 0$ then $M=\frac{\operatorname{tr}(H)}{2}{\mathtt{I}}-H$ has eigenvalues $\lambda_{1}\geq\lambda_{2}\geq|\lambda_{3}|$.
@@
If $\eta_{1}\geq\eta_{2}\geq\eta_{3}\geq 0$ are eigenvalues of $H$ then $\frac{\operatorname{tr}(H)}{2}-\eta_{i}$, $i=1,2,3$ are the eigenvalues of $M$. Sorted in decreasing order we get $\lambda_{1}=\frac{1}{2}(\eta_{1}+\eta_{2}-\eta_{3})$, $\lambda_{2}=\frac{1}{2}(\eta_{1}-\eta_{2}+\eta_{3})$ and $\lambda_{3}=\frac{1}{2}(-\eta_{1}+\eta_{2}+\eta_{3})$. Only $\lambda_{3}$ can be negative and clearly $\lambda_{2}\pm\lambda_{3}\geq 0$. ∎
The case $\lambda_{3}<0$ occurs when the leading eigenvalue of $H$ is significantly larger than the other two, i.e. the estimate $\bm{\omega}$ is more certain in a specific direction. This is actually a rather common scenario in image-based estimation as the in-plane part of the rotation is usually better constrained than the out-of-plane rotation angles.
@@
Incorporating the anisotropic objective (11) in the standard relaxation (SDP-O(3)-iso) may seem like a straightforward extension. We refer to this approach as SDP-O(3)-aniso. The source of the problem with this approach is that it ignores the determinant constraint and optimizes over $O(3)$. To illustrate this problem, we ran $1000$ synthetic anisotropic problem instances (the data generation protocol follows Section 5.1 with the eigenvalues of inverse Hessians sampled from $[0.1,1]$). Fig. 2 shows that the standard rotation averaging relaxation is never able to recover a rank-$3$ solution [^3] To determine rank, we used the smallest number $N$ of singular values (i.e., the first $N$ largest values) that sum up to $>99.9\%$ of the total sum. for any of the problem instances. In contrast, our relaxation SDP-cSO(3), presented further in Section 4.3, returns rank-$3$ solutions in all cases.
Figure 2: The results of running 1000 synthetic instances of rotation averaging with an anisotropic objective. In contrast to the proposed approach the standard convex relaxations give no solutions of rank $3$.
To better understand the problem, we consider terms of type $-\langle M\tilde{R},R\rangle$ that constitute the objective function. As we have seen in Section 4.1, these are locally accurate and allow uncertainties of $\tilde{R}$ to be propagated to the rotation averaging stage. However, for successful global optimization one has to guarantee that there is no other matrix in the feasible set that gives a lower cost than $\tilde{R}$. The following theorem shows, that when viewed as a function over $O(3)$, $\tilde{R}$ does typically not result in the smallest objective value:
@@
We minimize $-\langle M\tilde{R},R\rangle$ by computing KKT-points over $O(3)$. Introducing a (symmetric) Lagrange multiplier $\Lambda$ for the constraint $RR^{\top}-{\mathtt{I}}=0$ and differentiating
@@
$-\langle UDU^{\top}\tilde{R},USU^{\top}\tilde{R}\rangle=\pm\lambda_{1}\pm\lambda_{2}\pm\lambda_{3},$ (15)
where $\lambda_{1},\lambda_{2},\lambda_{3}$ are the eigenvalues of $M$. According to Lemma 1 the smallest eigenvalue $\lambda_{3}$ of $M$ is negative if $M$ is indefinite. In this case the objective function is minimized over $O(3)$ by choosing $S=\operatorname{diag}(1,1,-1)$ which yields the solution $R=USU^{\top}\tilde{R}$ and $-\langle M\tilde{R},R\rangle=-\lambda_{1}-\lambda_{2}-|\lambda_{3}|$. In contrast $-\langle M\tilde{R},\tilde{R}\rangle=-\lambda_{1}-\lambda_{2}+|\lambda_{3}|$, showing that the smallest value of $f$ is $f(R)=-2|\lambda_{3}|$.
Since $SO(3)\subset O(3)$ the KKT points of $SO(3)$ are those which satisfy $\det(R)=1$, which implies that the matrix $S$ has to have either zero or two elements that are $-1$. Therefore the above choice of $S$ is infeasible, and the optimal $S$ is instead given by $S=I$, which shows that the minimum value of $f$ is $f(\tilde{R})=0$. ∎
The above result shows that loss-terms of the type $-\langle M\tilde{R},R\rangle$ will favor incorrect solutions (that can be far from the estimation $\tilde{R}$) when optimizing over $O(3)$, which explains the results in Figure 2. While strictly enforcing $R\in SO(3)$ is undesirable, as it results in non-convex constraints, the following result shows that it is enough to use $R\in\operatorname{conv}(SO(3))$ to resolve this issue.
@@
For a linear objective function it is well known that the optimal value is attained in an extreme point. For any set $C$, the extreme points of $\operatorname{conv}(C)$ are always contained in $C$ [34]. Further, since for every point $Q\in SO(3)$ the linear function $-\langle Q,R\rangle$ has its unique minimum in $R=Q$, it can be deduced that the set of extreme points of $\operatorname{conv}(SO(3))$ is exactly $SO(3)$. Thus, the minimum value of $f$ is the same over $SO(3)$ as over $\operatorname{conv}(SO(3))$. ∎

### v2

@@ Abstract
Rotation averaging is a key subproblem in applications of computer vision and robotics. Many methods for solving this problem exist, and there are also several theoretical results analyzing difficulty and optimality. However, one aspect that most of these have in common is a focus on the isotropic setting, where the intrinsic uncertainties in the measurements are not fully incorporated into the resulting optimization task. Recent empirical results suggest that moving to an anisotropic framework, where these uncertainties are explicitly included, can result in an improvement of solution quality. However, global optimization for rotation averaging has remained a challenge in this scenario. In this work we show how anisotropic costs can be incorporated in certifiably optimal rotation averaging. We also demonstrate how existing solvers, designed for isotropic situations, fail in the anisotropic setting. Finally, we propose a stronger relaxation and empirically show that it recovers global optima in all tested datasets and leads to more accurate reconstructions in almost all scenes. [^1] This work was supported by the Wallenberg Artificial Intelligence, Autonomous Systems and Software Program (WASP), funded by the Knut and Alice Wallenberg Foundation, and the Swedish Research Council (#2023-05341).
@@ 1 Introduction
In large part this is because of the great importance it has for many problems within computer vision and robotics. Traditionally, it has been a core component of non-sequential structure from motion (SfM), see e.g. [29, 30, 35]. The key strength here is that rotation averaging optimizes the orientations of all the cameras at the same time, allowing recovery methods to avoid incrementally increasing errors (drift) [14] that easily occur in sequential SfM methods. More recently, [7, 9, 10] point out the importance of rotation averaging to improve the accuracy and speed of simultaneous localization and mapping (SLAM).
@@
Further, we empirically verify—on a number of synthetic and real datasets—that this new relaxation is sufficient to recover the globally optimal solution, and that our anisotropic model generally achieves solutions that are more accurate than the standard isotropic chordal penalty. In summary, our main contributions are:
@@ 2 Related Work
The works [3, 23] also consider certifiable optimization of the anisotropic chordal distance. However, their SDP relaxation is based on the Cayley mapping and the relation to our formulation is somewhat unclear. In addition, [3, 23] have to introduce redundant constraints in order to get a “reasonably tight” SDP relaxation. It is likely, in the light of our analysis in this paper, and the discussion in [6] that this is a result of the authors only enforcing $O(3)$ membership for their optimization.
It is also worth mentioning that—while this paper looks at the anisotropic $L_{2}$ chordal distance—it is possible to consider other distances between rotations as well as their robust counterparts. E.g., [47, 48] consider the geodesic distance. Further, in [11, 43, 46] the authors analyze and look at different (robustified) distances in order to obtain a “robust” formulation (resulting in the solution being less susceptible to outliers). Outliers are handled via the introduction of auxiliary variables in the SDP formulation in [38, 8].
From an algorithmic point of view, the SDP relaxation can be solved by general-purpose solvers, e.g. [1]. There are also several dedicated solvers, constructed specifically for the rotation averaging problem. [18, 36] use coordinate descent (fixing all but one row/column and minimizing), [28] use a primal-dual update rule, and in [12, 49] the authors combine a global coordinate descent approach with a local step. An early approach [44] looks at eigenvectors of a specific matrix. A particularly powerful method used in [5, 16] is that of the Riemannian staircase, where one optimizes using block matrices of a certain dimension and increases (“climb the staircase”) the dimension until a global optimum is found. The authors show that such a method can be made highly effective.
Like most work on rotation averaging, we focus on point estimates and assume unimodal Gaussian noise in the observations. If the input data has multiple modes (e.g. due to perceptual aliasing), the particle-based algorithm proposed in [4] provides more expressive marginals. Perceptual aliasing can also be explicitly addressed by leveraging cycle consistency constraints (e.g. [50, 25, 42]).
@@ 3 Certifiably Optimal Rotation Averaging
The derivation of a strong convex relaxation relies on two observations: Firstly, since the set of rotation matrices have constant (Frobenius) norm, the objective function can be replaced by $\sum_{i\neq j}\langle\tilde{R}_{ij},R_{j}R_{i}^{\top}\rangle$. If we let $\mathbf{R}=\begin{bmatrix}R_{1}^{\top}&R_{2}^{\top}&\ldots\end{bmatrix}^{\top}$ and $\mathbf{\tilde{R}}$ be the matrix containing the blocks $\tilde{R}_{ij}$ (and zeros where no relative rotation estimate is available) we can simplify the objective to $\langle\mathbf{\tilde{R}},\mathbf{R}\mathbf{R}^{\top}\rangle$, which is linear in the relative rotations. Secondly, by orthogonality we have $R_{i}R_{i}^{\top}={\mathtt{I}}$, resulting in the Quadratically Constrained Quadratic Program (QCQP)
@@
$d_{2}^{*}:=\min_{\mathbf{X}\succeq 0}$ $\,-\langle\mathbf{\tilde{R}},\mathbf{X}\rangle\quad\text{s.t. }\forall i:X_{ii}={\mathtt{I}}.$ (SDP-O(3)-iso)
The above program is convex and can therefore be solved reliably using general purpose [1, 26] or specialized solvers [16, 5, 18]. The difference $p^{*}-d_{2}^{*}$ between objective values is referred to as the duality gap. It is clear that we can construct $\mathbf{X}=\mathbf{R}\mathbf{R}^{\top}$ from a solution $\mathbf{R}$ of (2), which is feasible in (SDP-O(3)-iso) and therefore $p^{*}\geq d_{2}^{*}$. Moreover, if the solution $\mathbf{X}$ to (SDP-O(3)-iso) has $\operatorname{rank}(\mathbf{X})=3$ then it can be factorized into $\mathbf{X}=\mathbf{R}\mathbf{R}^{\top}$, with $\mathbf{R}$ being feasible in (2), which yields $p^{*}=d_{2}^{*}$ and certifies that $\mathbf{R}$ is globally optimal in (2). Thus, a solution of the original non-convex QCQP is obtained via a convex SDP program.
@@ 4 Anisotropic Rotation Averaging
While the use of the chordal distance often results in tight SDP relaxations, a downside is that it does not take into account the uncertainty of the relative rotation estimates $\tilde{R}_{ij}$. For two-view relative pose problems, it is frequently the case that the reprojection errors are much less affected by rotations in certain directions than others. In Figure 1 we illustrate this on a real two-view problem. The left image shows the ground truth solution (obtained after bundle adjustment). To generate the graphs to the right, we sampled rotations of the second camera around its x-,y-, and z-axes between 5-5 and $5$ degrees. For each sampled rotation, we found the best 3D point locations and position of the second camera, and computed the sum-of-squared (calibrated) reprojection errors (solid curves). The result shows that rotations around the black/upward axis affects the reprojection errors significantly less than the others, resulting in a direction of larger uncertainty. By performing sensitivity analysis (i.e. by looking at the singular values of the Jacobian) one can arrive at the similar conclusion. It is therefore desirable to propagate this information to the rotation averaging stage to increase the accuracy of the model. The dashed curves (that are almost indistinguishable from the solid ones) show the approximations, presented below, that can be used for this purpose. In the supplementary material, we also show how isotropic rotation averaging gets negatively affected by the single noisy relative rotation (while the proposed method presented further remains unaffected).
@@ 4.1 Incorporating uncertainties
Consider a relative rotation $\tilde{R}_{ij}$ between cameras $i$ and $j$, that has been fully optimized using e.g. a Gauss-Newton method for two-view optimization. Using the exponential map we write $R_{ij}=e^{[\Delta\bm{\omega}_{ij}]_{\times}}\tilde{R}_{ij}$ [^2] Other parameterizations are discussed in the supplementary material., where $\Delta\bm{\omega}_{ij}$ is an axis-angle representation of the deviation from $\tilde{R}_{ij}$, i.e. the axis-angle of $R_{ij}\tilde{R}_{ij}^{\top}$. Since $\tilde{R}_{ij}$ is a minimum, gradient terms vanish and the two-view objective can locally be approximated by a quadratic function (up to an irrelevant constant)
$\tfrac{1}{2}\Delta\bm{\omega}_{ij}^{\top}H_{ij}\Delta\bm{\omega}_{ij}.$ (3)
We assume that $H_{ij}$ is positive semidefinite—usually given by a Gauss-Newton approximation $H_{ij}=J_{ij}^{\top}J_{ij}$, where $J_{ij}$ is the Jacobian of the underlying residual function. We remark that the two-view objective may depend on additional parameters such as camera positions and 3D point locations. Here we assume that these “nuisance” parameters have been marginalized out (using the Schur complement), and therefore we work with a reduced objective solely in terms of the rotation parameters $\Delta\bm{\omega}_{ij}$.
We assume Laplace’s approximation to be valid at least locally in order to quantify the uncertainty of $\Delta\bm{\omega}_{ij}$, i.e. identify the precision matrix with the Hessian $H_{ij}$. While in general the uncertainty of a $3\times 3$ matrix $R_{ij}$ is represented by a $9\times 9$ covariance/precision matrix, we aim for a linear cost (in terms of rotation matrices) in order to achieve a strong relaxation, which means that the objective is restricted to an inner product between $R_{ij}$ and a constant but input-dependent matrix. We first determine a matrix $M_{ij}$ such that
$\begin{split}\Delta\bm{\omega}_{ij}^{\top}H_{ij}\Delta\bm{\omega}_{ij}&=\operatorname{tr}\left([\Delta\bm{\omega}_{ij}]_{\times}^{\top}M_{ij}[\Delta\bm{\omega}_{ij}]_{\times}\right)\\ &=-\operatorname{tr}\left(M_{ij}[\Delta\bm{\omega}_{ij}]_{\times}^{2}\right).\end{split}$ (4)
Using the identity $\mathbf{v}]_{\times}^{2}=\mathbf{v}^{\top}\mathbf{v}{\mathtt{I}}-\mathbf{v}\mathbf{v}^{\top}$, we obtain that $H_{ij}=\operatorname{tr}(M_{ij}){\mathtt{I}}-M_{ij}$. Taking the trace on both sides further shows that $\operatorname{tr}(H_{ij})=2\operatorname{tr}(M_{ij})$, resulting in
$M_{ij}=\tfrac{\operatorname{tr}(H_{ij})}{2}{\mathtt{I}}-H_{ij}.$ (5)
@@
$\Delta\bm{\omega}_{ij}]_{\times}\approx R_{ij}\tilde{R}_{ij}^{\top}-{\mathtt{I}},$ (6)
@@
$\tfrac{1}{2}\operatorname{tr}\big{(}(R_{ij}\tilde{R}_{ij}^{\top}-{\mathtt{I}})^{\top}M_{ij}(R_{ij}\tilde{R}_{ij}^{\top}-{\mathtt{I}})\big{)}.$ (7)
@@
$\operatorname{tr}\big{(}M_{ij}\tilde{R}_{ij}R_{ij}^{\top}\big{)}=-\langle M_{ij}\tilde{R}_{ij},R_{ij}\rangle.$ (8)
We observe that this objective can be viewed as a negative log-likelihood of a Langevin distribution $\mathcal{L}(R_{ij},M_{ij})$, e.g. [13], with a density function of the type
$p(\tilde{R}_{ij})\propto e^{\langle M_{ij}\tilde{R}_{ij},R_{ij}\rangle},$ (9)
and that generating rotations $\tilde{R}_{ij}$ according to (9) is approximately (up to second order terms) equivalent to generating axis-angle vectors $\Delta\bm{\omega}_{ij}$ from a normal distribution $\mathcal{N}(0,H^{-1}_{ij})$ with density
$p(\Delta\bm{\omega}_{ij})\propto e^{-\frac{1}{2}\Delta\bm{\omega}_{ij}^{\top}H_{ij}\Delta\bm{\omega}_{ij}}.$ (10)
Hence, the above derivations suggest to generalize the rotation averaging approach by replacing the relative estimates $\tilde{R}_{ij}$ with a weighted version $M_{ij}\tilde{R}_{ij}$, where $M_{ij}$ is based on the Hessian $H_{ij}$ of the relative pose problem as outlined above. Therefore, our proposed objective to minimize is
$\sum\nolimits_{i\neq j}\langle M_{ij}\tilde{R}_{ij},R_{j}R_{i}^{\top}\rangle=-\langle\mathbf{N},\mathbf{R}\mathbf{R}^{\top}\rangle,$ (11)
where $\mathbf{N}$ is a symmetric block matrix containing the blocks $(M_{ij}\tilde{R}_{ij})^{\top}$ for $i<j$, 0 when $i=j$, and $M_{ji}\tilde{R}_{ji}$ for $i>j$.
We conclude this section by remarking that—in contrast to $H_{ij}$—the matrix $M_{ij}$ will generally not be positive semidefinite as the following lemma shows.
@@
If $H_{ij}\succeq 0$ then $M_{ij}=\frac{\operatorname{tr}(H_{ij})}{2}{\mathtt{I}}-H_{ij}$ has eigenvalues $\lambda_{1}\geq\lambda_{2}\geq|\lambda_{3}|$.
@@
If $\eta_{1}\geq\eta_{2}\geq\eta_{3}\geq 0$ are eigenvalues of $H_{ij}$ then $\frac{\operatorname{tr}(H_{ij})}{2}-\eta_{i}$, $i=1,2,3$ are the eigenvalues of $M_{ij}$. Sorted in decreasing order we get $\lambda_{1}=\frac{1}{2}(\eta_{1}+\eta_{2}-\eta_{3})$, $\lambda_{2}=\frac{1}{2}(\eta_{1}-\eta_{2}+\eta_{3})$ and $\lambda_{3}=\frac{1}{2}(-\eta_{1}+\eta_{2}+\eta_{3})$. Only $\lambda_{3}$ can be negative and clearly $\lambda_{2}\pm\lambda_{3}\geq 0$. ∎
The case $\lambda_{3}<0$ occurs when the leading eigenvalue of $H_{ij}$ is significantly larger than the other two, i.e. the estimate $\bm{\omega}_{ij}$ is more certain in a specific direction. This is actually a rather common scenario in image-based estimation as the in-plane part of the rotation is usually better constrained than the out-of-plane rotation angles. Table 1 shows the percentage of indefinite matrices $M_{ij}$ for the real datasets tested in Section 5.2.
@@
Incorporating the anisotropic objective (11) in the standard relaxation (SDP-O(3)-iso) may seem like a straightforward extension. We refer to this approach as SDP-O(3)-aniso. The source of the problem with this approach is that it ignores the determinant constraint and optimizes over $O(3)$. To illustrate this problem, we ran $1000$ synthetic anisotropic problem instances (the data generation protocol follows Section 5.1 with the eigenvalues of inverse Hessians sampled from $[0.1,1]$). Fig. 2 shows that the standard rotation averaging relaxation is never able to recover a rank-$3$ solution [^4] To determine rank, we used the smallest number $N$ of singular values (i.e., the first $N$ largest values) that sum up to $99.9\%$ of the total sum. for any of the problem instances. In contrast, our relaxation SDP-cSO(3), presented further in Section 4.3, returns rank-$3$ solutions in all cases.
Figure 2: The results of running 1000 synthetic instances of rotation averaging with an anisotropic objective. In contrast to the proposed approach the standard convex relaxations give no solutions of rank 33.
To better understand the problem, we consider terms of type $\langle M\tilde{R},R\rangle$ [^5] For the remainder of this section we focus on a single relative rotation and therefore omit the subscripts. that constitute the objective function. As we have seen in Section 4.1, these are locally accurate and allow uncertainties of $\tilde{R}$ to be propagated to the rotation averaging stage. However, for successful global optimization one has to guarantee that there is no other matrix in the feasible set that gives a lower cost than $\tilde{R}$. The following theorem shows, that when viewed as a function over $O(3)$, $\tilde{R}$ does typically not result in the smallest objective value:
@@
We minimize $\langle M\tilde{R},R\rangle$ by computing KKT-points over $O(3)$. Introducing a (symmetric) Lagrange multiplier $\Lambda$ for the constraint $RR^{\top}-{\mathtt{I}}=0$ and differentiating
@@
$\langle UDU^{\top}\tilde{R},USU^{\top}\tilde{R}\rangle=\pm\lambda_{1}\pm\lambda_{2}\pm\lambda_{3},$ (15)
where $\lambda_{1},\lambda_{2},\lambda_{3}$ are the eigenvalues of $M$. According to Lemma 1 the smallest eigenvalue $\lambda_{3}$ of $M$ is negative if $M$ is indefinite. In this case the objective function is minimized over $O(3)$ by choosing $S=\operatorname{diag}(1,1,-1)$ which yields the solution $R=USU^{\top}\tilde{R}$ and $\langle M\tilde{R},R\rangle=-\lambda_{1}-\lambda_{2}-|\lambda_{3}|$. In contrast $\langle M\tilde{R},\tilde{R}\rangle=-\lambda_{1}-\lambda_{2}+|\lambda_{3}|$, showing that the smallest value of $f$ is $f(R)=-2|\lambda_{3}|$.
Since $SO(3)\subset O(3)$ the KKT points of $SO(3)$ are those which satisfy $\det(R)=1$, which implies that the matrix $S$ has to have either zero or two elements that are 1-1. Therefore the above choice of $S$ is infeasible, and the optimal $S$ is instead given by $S=I$, which shows that the minimum value of $f$ is $f(\tilde{R})=0$. ∎
The above result shows that loss-terms of the type $\langle M\tilde{R},R\rangle$ will favor incorrect solutions (that can be far from the estimation $\tilde{R}$) when optimizing over $O(3)$, which explains the results in Figure 2. While strictly enforcing $R\in SO(3)$ is undesirable, as it results in non-convex constraints, the following result shows that it is enough to use $R\in\operatorname{conv}(SO(3))$ to resolve this issue.
@@
For a linear objective function it is well known that the optimal value is attained in an extreme point. For any set $C$, the extreme points of $\operatorname{conv}(C)$ are always contained in $C$ [37]. Further, since for every point $Q\in SO(3)$ the linear function $\langle Q,R\rangle$ has its unique minimum in $R=Q$, it can be deduced that the set of extreme points of $\operatorname{conv}(SO(3))$ is exactly $SO(3)$. Thus, the minimum value of $f$ is the same over $SO(3)$ as over $\operatorname{conv}(SO(3))$. ∎